an employer has 5 female and 7 male qualified for 5 vacant positions. In how many ways can he hire 2 males and 3 females?
Answer by stanbon(75887)

an employer has 5 female and 7 male qualified for 5 vacant positions. In how many ways can he hire 2 males and 3 females?
-----
Ans: 5C3*7C2 = 10*21 = 210 ways
------------
Cheers,
Stan H.
=============